Determine the general solution of y' + [1-3x^2/x(1-x^2)]y = 1/1-x^2
1
Expert's answer
2020-05-12T17:19:37-0400

Given

"y' + \frac{1-3x^2}{x(1-x^2)} y =\frac{1} {1-x^2}"

This is a linear equation with

"p(x)=\frac{1-3x^2}{x(1-x^2)},\ \ \ q(x) =\frac{1} {1-x^2}"

Since

"\begin{aligned}
\frac{1-3x^2}{x(1-x^2)}&=\frac{A}{x }+\frac{B}{1-x}+\frac{C}{1+x}\\
&= \frac{A (1-x^2)+Bx(1+x)+Cx(1-x)}{x(1-x^2)}\\
1-3x^2&=A (1-x^2)+Bx(1+x)+Cx(1-x)
\end{aligned}"

Then

"\begin{aligned}
\text{At } \ x&=0\ \ \ \to \ A=1\\
 \text{At } \ x&=-1\ \ \ \to \ C=1\\
 \text{At } \ x&= 1\ \ \ \to \ B=-1
\end{aligned}"

Hence

"\begin{aligned}
\int p(x)dx&=\int \frac{1-3x^2}{x(1-x^2)}dx\\
&=\int \frac{dx}{x }-\int \frac{dx}{1-x}+\int \frac{dx}{1+x}\\
&=\ln x+\ln (1-x)+\ln (1+x)\\
&=\ln x(1-x^2)
\end{aligned}"

Since the integral factor given by

"\begin{aligned}
\mu&=e^{\int p(x)dx}\\
&=e^{\ln x(1-x^2)}\\
&= x(1-x^2)
\end{aligned}"

Then the solution is

"\begin{aligned}
\mu y&=\int \mu q(x)dx\\
x(1-x^2) y &= \int x(1-x^2)\frac{1}{1-x^2}dx\\
&=\int xdx\\
&= \frac{1}{2}x^2+C
\end{aligned}"

Hence

"y= \frac{ x^2}{2x(1-x^2)}+\frac{ C }{2x(1-x^2)}"